OpenStudy (anonymous):

What is the equation of a line that has a slope of 5/6 and passes through the point (-5,-6)? PLEASE HELP!

OpenStudy (anonymous):

Use the equation of a line y=mx+b. Let m=5/6, sbstitute into equation with te point given. -6=(5/6)-5 + b. Solve for b, to give y= 5/6x-11/6

OpenStudy (anonymous):

ok so if the slope was perpendicular would the answer be y=-6/5x-12